OpenStudy (perl):

eliminate arbitrary constants using determinant. y=c1*exp(x) + c2*exp(2x). so they want us to find a 2nd order diff. equation

OpenStudy (perl):

y' = c1*exp(x) + 2*c2*exp(2x) y'' = c1*exp(x) + 4*c2*exp(2x)